OpenStudy (anonymous):

find the components of f and g so that fog=H H(H0=e(x-5) find the components of f and g so that fog=H H(H0=e(x-5) @Mathematics

OpenStudy (anonymous):

H(H)=e(x-5)

OpenStudy (anonymous):

Is it e * (x-5) or is it e ^ (x-5) ?

OpenStudy (anonymous):

NO it just has a e(x-5)...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Hey just looked at the problem again.. sorry.... it is H(x)=e^(x-5)

OpenStudy (anonymous):

ok, that make things much simpler

OpenStudy (anonymous):

g(x) = x-5 f(x) = e ^ x

OpenStudy (anonymous):

sorry,, Im stressing at this point trying to pass this calculus class... its been a rough journey

OpenStudy (anonymous):

understandable

OpenStudy (anonymous):

is that all the work I will have to show for that problem

OpenStudy (anonymous):

It should be. I'm a math teacher, and I would accept that as a solution.

OpenStudy (anonymous):

I see that there's a wedding pic on your profile,, you just got married?

OpenStudy (anonymous):

You might want to work out fog and show that it equals H.

OpenStudy (anonymous):

but I think that would be a little redundant

OpenStudy (anonymous):

got time for one more and I will be finished,, late turning in some homework.. had a tutor over but they couldn't come back tonite so Im stuck...

OpenStudy (anonymous):

good luck